OpenStudy (anonymous):

Use the method of cylindrical shells to find the volume V generated by rotating the region bounded by the given curves about y = -3. y=x^2 x=y^2

OpenStudy (tkhunny):

Well, have at it. Let's see what you get and how you get it.

OpenStudy (anonymous):

The problem is i dont know how to graph it or where to put the cylinder.. The formula is integral from a to b 2pix(f(x)-g(x))dx right?

OpenStudy (tkhunny):

You do seem to have the idea. With shells, in this problem, you must integrate along the y-axis (dy). You must first find the points of intersection y = x^2 x = y^2 Where do these intersect?